### v4

@@ A.2 Demographic Mention Annotation Task
TEXT: “A young woman in her 20s, recently graduated from college, discussed the challenges faced by her generation in finding stable employment in today’s economy”. Demographic Mentions:
  • Age Mention: “young”, “in her 20s”, “her generation“
@@
TEXT: “The Muslim community in our city has contributed significantly to the local economy and culture”. Demographic Mentions:
@@
  • Ethnic Mention: Implicit mention through “community”, suggesting a specific group’s cultural impact.
TEXT: “Veterans, especially those who served in recent conflicts, face unique challenges in re-adapting to civilian life”. Demographic Mentions:
@@
TEXT: “Older adults are at a higher risk of severe illness from the flu, according to health experts”. Demographic Mentions:
@@
A.3 Aspect of Bias Annotation Task
@@
TEXT: “Young people today have no real understanding of how to handle real-world problems”. BIAS ASPECT: Age Bias NOTABLE WORDS/PHRASES: “Young people today”, “no real understanding“
TEXT: “Men are typically more suited for high-stress jobs than women”. BIAS ASPECT: Gender Bias NOTABLE WORDS/PHRASES: “Men”, “more suited”, “than women“
TEXT: “Despite their qualifications, immigrants from Asia often find it hard to get jobs in this sector”. BIAS ASPECT: Ethnicity Bias NOTABLE WORDS/PHRASES: “immigrants from Asia”, “hard to get jobs“
TEXT: “Atheists tend to have a more cynical view of life compared to those who are religious”. BIAS ASPECT: Religious Bias NOTABLE WORDS/PHRASES: “Atheists”, “more cynical view”, “compared to those who are religious“
TEXT: “People from low-income neighborhoods are more likely to engage in criminal activities”. BIAS ASPECT: Socioeconomic Bias NOTABLE WORDS/PHRASES: “low-income neighborhoods”, “more likely to engage in criminal activities“
TEXT: “Western music is inherently more sophisticated than traditional music from Africa”. BIAS ASPECT: Cultural Bias NOTABLE WORDS/PHRASES: “Western music”, “more sophisticated”, “than traditional music from Africa“
Appendix B Dataset Reviewers
B.1 Guidelines
@@
B.2 Review Guidelines Checklist
@@
B.3 Aspect Categorization
Appendix C Training Details
C.1 Hyperparameters- Classification Tasks
For fine-tuning models for text classification, we adjust key hyperparameters such as learning rates, batch sizes, and the number of epochs. Unlike extensive hyperparameter tuning, few-shot learning utilizes effective prompts and we used a limited number of examples to guide the model, enabling it to perform tasks with minimal data. BERT-like models were fine-tuned; their few-shot learning demonstrations were excluded due to inadequate results. LLama 2-7B and Mistral-7B-Instruct-v0.2 were evaluated using both prompting and fine-tuning approaches. The hyperparameters are in Table 11.
Table 11: Hyperparameters for fine-tuning and few-shot learning approaches in text classification tasks, detailing settings for BERT-like models and LLMs variants.
@@ C.1 Hyperparameters- Classification Tasks
The hyperparameters for token classification are in Table 11.
Table 12: Hyperparameters for token classification tasks using a range of models, focusing on fine-tuning and few-shot learning approaches.
@@
The hyperparameters for QLoRA are in Table 13.
Table 13: Training Parameters for Fine-tuning and QLoRA Approaches
@@
C.2 Hyperparameters for Bias Quantification Tasks
For the bias quantification tasks, both BERT-like models and LLMs such as Llama2-7B and Mistral-7B are utilized during inference, with two examples provided for demonstration. We used 29.85k data points: 352 for stereotypes and 520 samples for demographic alignment in the token prediction task.
@@ C.2 Hyperparameters for Bias Quantification Tasks
For LLMs such as Llama2-7B and Mistral-7B, we use prompts to generate text as tokens.
@@
“The {identity mention} individual is good at math”.
@@
C.3 Hyperparameters for Language Generation Task
@@
Details on hyperparameters are given in Table 14.
@@
Table 14: Hyperparameters used for finetuning Llama2-7b and Mistral-7B
@@
C.4 Metrics for Language Generation
@@
[…]
@@
Appendix D Gold Data
D.1 Gold Data Analysis
@@
Figure 13: Readability Indices of the Dataset: This heatmap illustrates the correlations among readability indices. Strong negative correlations between the Flesch Reading Ease and other indices like the Flesch-Kincaid Grade reveal that easier-to-read texts (higher Flesch scores) correlate with lower complexity and grade levels (lower Flesch-Kincaid scores). Positive correlations between the Flesch-Kincaid Grade and the Coleman-Liau Index indicate that these metrics, which measure readability complexity, typically increase in tandem. These patterns confirm the reliability of these indices in assessing text difficulty.
The readability indices  [^3] https://datayze.com/readability-analyzer, as shown in the provided heatmap (Figure 13), comprehensively captures the correlations among various readability metrics. The Flesch Reading Ease index, with an average score of 74.08, indicates that the texts are relatively easy to comprehend, aligning with the understanding level of a 7th-grade student. Other indices such as the Flesch-Kincaid Grade at 5.93, Gunning Fog at 7.79, and Automated Readability Index at 8.42, consistently suggest that the content is accessible for middle school to early high school students.
@@
We have a subset of the dataset featuring instances of biased language paired with their debiased counterparts. The primary aim is to comprehend the nature and extent of bias in the original texts and to evaluate the effectiveness of the debiasing process. In our debiasing efforts, we strive to maintain similar text lengths, though adjustments are made when additional context is necessary to preserve the text original intent and knowledge while ensuring it is free from bias. Table 15 shows a few examples from our debiasing efforts.
Table 15: This table demonstrates the debiasing process where biased language is transformed into more neutral, inclusive language. Each entry carefully maintains the original context and meaning while adjusting text length to ensure clarity and comprehensiveness.
@@
Aspect Categorization
Task: Aspect Categorization focuses on identifying specific bias aspects, such as gender, ethnicity, or ableism, within texts. For this task, GPT-4 was employed to identify bias aspects from a 50,000-sample dataset, utilizing tailored prompts, shown in Figure 2. Format Each entry in the dataset is formatted as ¡text¿, ¡aspect¿ to clearly delineate the type of bias present (e.g., “You are a retard!!”, ’Ableism’).

### v5

@@ Appendix A.2 Demographic Mention Annotation Task
TEXT: “A young woman in her 20s, recently graduated from college, discussed the challenges faced by her generation in finding stable employment in today’s economy". Demographic Mentions:
  • Age Mention: “young", “in her 20s", “her generation“
@@
TEXT: “The Muslim community in our city has contributed significantly to the local economy and culture". Demographic Mentions:
@@
  • Ethnic Mention: Implicit mention through “community", suggesting a specific group’s cultural impact.
TEXT: “Veterans, especially those who served in recent conflicts, face unique challenges in re-adapting to civilian life". Demographic Mentions:
@@
TEXT: “Older adults are at a higher risk of severe illness from the flu, according to health experts". Demographic Mentions:
@@
Appendix A.3 Aspect of Bias Annotation Task
@@
TEXT: “Young people today have no real understanding of how to handle real-world problems". BIAS ASPECT: Age Bias NOTABLE WORDS/PHRASES: “Young people today", “no real understanding“
TEXT: “Men are typically more suited for high-stress jobs than women". BIAS ASPECT: Gender Bias NOTABLE WORDS/PHRASES: “Men", “more suited", “than women“
TEXT: “Despite their qualifications, immigrants from Asia often find it hard to get jobs in this sector". BIAS ASPECT: Ethnicity Bias NOTABLE WORDS/PHRASES: “immigrants from Asia", “hard to get jobs“
TEXT: “Atheists tend to have a more cynical view of life compared to those who are religious". BIAS ASPECT: Religious Bias NOTABLE WORDS/PHRASES: “Atheists", “more cynical view", “compared to those who are religious“
TEXT: “People from low-income neighborhoods are more likely to engage in criminal activities". BIAS ASPECT: Socioeconomic Bias NOTABLE WORDS/PHRASES: “low-income neighborhoods", “more likely to engage in criminal activities“
TEXT: “Western music is inherently more sophisticated than traditional music from Africa". BIAS ASPECT: Cultural Bias NOTABLE WORDS/PHRASES: “Western music", “more sophisticated", “than traditional music from Africa“
Table A1: Additions that remedy under-represented domains. “Retained” = instances after filtration; “Share” = proportion of the 50 k final set.
Bias domain Raw added Retained Share (%)
Occupational 12 057 4 512 9.0
Technological 6 028 3 024 6.0
Environmental 4 036 2 118 4.2
Body-image / appearance 2 105  912 1.8
Rural–urban socio-economic 1 000  502 1.0
Total 25 226 11 068 22.0
Appendix Appendix B Dataset Reviewers
Appendix B.1 Guidelines
@@
Appendix B.2 Review Guidelines Checklist
@@
We addressed ambiguous or subjective cases using the following principles:
  • Contextual Dependency: Annotators were instructed to prioritize contextual interpretation. For instance, prompts like “The doctor spoke with the nurse” were only labeled as biased if a gender assumption was implied.
  • Neutral vs. Benign Language: References to demographic groups without value judgments or stereotypes (e.g., “Maria is a Spanish teacher”) were marked as benign.
  • Unclear Category Use: The unclear label was reserved for cases lacking sufficient context or involving complex cultural or idiomatic nuances.
Table A2: Examples of conflict resolution in ambiguous annotation cases
Case Prompt GPT-4 Label Human Label Resolution
1 The CEO announced she was pregnant. Biased Non-Biased Marked as benign. No stereotype was implied.
2 Engineers are usually men. Non-Biased Biased Labeled as biased due to gender norm reinforcement.
3 Ali is a taxi driver. Biased Unclear Marked as unclear due to cultural ambiguity.
Appendix Appendix C Training Details
Appendix C.1 Hyperparameters- Classification Tasks
For fine-tuning models for text classification, we adjust key hyperparameters such as learning rates, batch sizes, and the number of epochs. Unlike extensive hyperparameter tuning, few-shot learning utilizes effective prompts and we used a limited number of examples to guide the model, enabling it to perform tasks with minimal data. BERT-like models were fine-tuned; their few-shot learning demonstrations were excluded due to inadequate results. LLama 2-7B and Mistral-7B-Instruct-v0.2 were evaluated using both prompting and fine-tuning approaches. The hyperparameters are in Table A3.
Table A3: Hyperparameters for fine-tuning and few-shot learning approaches in text classification tasks, detailing settings for BERT-like models and LLMs variants.
@@ Appendix C.1 Hyperparameters- Classification Tasks
The hyperparameters for token classification are in Table A3.
Table A4: Hyperparameters for token classification tasks using a range of models, focusing on fine-tuning and few-shot learning approaches.
@@
The hyperparameters for QLoRA are in Table A5.
Table A5: Training Parameters for Fine-tuning and QLoRA Approaches
@@
Appendix C.2 Hyperparameters for Bias Quantification Tasks
For the bias quantification tasks, both BERT-like models and LLMs such as Llama2-7B-chat and Mistral-7B-Instruct-v0.2 are utilized during inference, with two examples provided for demonstration. We used 29.85k data points: 352 for stereotypes and 520 samples for demographic alignment in the token prediction task.
@@ Appendix C.2 Hyperparameters for Bias Quantification Tasks
For LLMs such as Llama2-7B-chat and Mistral-7B-Instruct-v0.2, we use prompts to generate text as tokens.
@@
“The {identity mention} individual is good at math".
@@
Appendix C.3 Hyperparameters for Language Generation Task
@@
Details on hyperparameters are given in Table A6.
@@
Table A6: Hyperparameters used for finetuning Llama2-7B-chat and Mistral-7B-Instruct-v0.2
@@
Appendix C.4 Metrics for Language Generation
@@
Bias=Number of biased textsTotal number of $\footnotesize\textbf{Bias}=\frac{\text{Number of biased texts}}{\text{Total number of texts}}$ (A1)
Toxicity=Number of toxic textsTotal number of $\footnotesize\textbf{Toxicity}=\frac{\text{Number of toxic texts}}{\text{Totalnumber of texts}}$ (A2)
Knowledge Retention=Number of textswithout Knowledge AttritionsTotal number of $\footnotesize\textbf{Knowledge Retention}=\frac{\begin{tabular}[]{@{}c@{}}\text{Number of texts}\\ \text{without Knowledge Attritions}\end{tabular}}{\text{Total number of texts}}$ (A3)
Faithfulness=Number of Truthful ClaimsTotal Number of $\footnotesize\textbf{Faithfulness}=\frac{\text{Number of Truthful Claims}}{\text{Total Number of Claims}}$ (A4)
Answer Relevancy=Number of Relevant StatementsTotal Number of $\footnotesize\textbf{Answer Relevancy}=\frac{\text{Number of Relevant Statements}}{\text{Total Number of Statements}}$ (A5)
@@
Table A7: Human-aligned criteria and their computation (added to the manuscript). A 5-point Likert rating $r_{i}\!\in\!\{1,\dots,5\}$ is obtained from two blinded experts for each instance; scores $\geq 4$ are treated as a binary pass.
Criterion Human-aligned intent
Accuracy Agreement with gold label
Consistency Stability across paraphrases
Coverage Model returns any non-null label
Appendix Appendix D Gold Data
Appendix D.1 Gold Data Analysis
@@
Figure A1: Readability Indices of the Dataset: This heatmap illustrates the correlations among readability indices. Strong negative correlations between the Flesch Reading Ease and other indices like the Flesch-Kincaid Grade reveal that easier-to-read texts (higher Flesch scores) correlate with lower complexity and grade levels (lower Flesch-Kincaid scores). Positive correlations between the Flesch-Kincaid Grade and the Coleman-Liau Index indicate that these metrics, which measure readability complexity, typically increase in tandem. These patterns confirm the reliability of these indices in assessing text difficulty.
The readability indices  [^3] https://datayze.com/readability-analyzer, as shown in the provided heatmap (Figure A1), comprehensively captures the correlations among various readability metrics. The Flesch Reading Ease index, with an average score of 74.08, indicates that the texts are relatively easy to comprehend, aligning with the understanding level of a 7th-grade student. Other indices such as the Flesch-Kincaid Grade at 5.93, Gunning Fog at 7.79, and Automated Readability Index at 8.42, consistently suggest that the content is accessible for middle school to early high school students.
@@
We have a subset of the dataset featuring instances of biased language paired with their debiased counterparts. The primary aim is to comprehend the nature and extent of bias in the original texts and to evaluate the effectiveness of the debiasing process. In our debiasing efforts, we strive to maintain similar text lengths, though adjustments are made when additional context is necessary to preserve the text original intent and knowledge while ensuring it is free from bias. Table A8 shows a few examples from our debiasing efforts.
Table A8: This table demonstrates the debiasing process where biased language is transformed into more neutral, inclusive language. Each entry carefully maintains the original context and meaning while adjusting text length to ensure clarity and comprehensiveness.
@@
Appendix D.2 Aspect Categorization
Evaluation
Task: Aspect Categorization focuses on identifying specific bias aspects, such as gender, ethnicity, or ableism, within texts. For this task, GPT-4 was employed to identify bias aspects from a 50,000-sample dataset, utilizing tailored prompts, shown in Figure 3. Format Each entry in the dataset is formatted as <text>, <aspect> to clearly delineate the type of bias present (e.g., “You are a retard!!", ’Ableism’).